40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Aug 5 13:28:43 2018, gautam, Update, CDS, c1lsc flaky 
    Reply  Mon Aug 6 00:26:21 2018, gautam, Update, CDS, More CDS woes CDSoverview.png
       Reply  Mon Aug 6 14:38:38 2018, gautam, Update, CDS, More CDS woes CDScrash.pngMC1failures.png
          Reply  Mon Aug 6 19:49:09 2018, gautam, Update, CDS, More CDS woes IMG_7106.JPG
             Reply  Tue Aug 7 11:30:46 2018, gautam, Update, CDS, More CDS woes 
                Reply  Tue Aug 7 22:28:23 2018, gautam, Update, CDS, More CDS woes 
                   Reply  Wed Aug 8 23:03:42 2018, gautam, Update, CDS, c1lsc model started 
                      Reply  Thu Aug 9 12:31:13 2018, gautam, Update, CDS, CDS status update 
                         Reply  Wed Aug 15 21:27:47 2018, gautam, Update, CDS, CDS status update 
                            Reply  Tue Sep 4 10:14:11 2018, gautam, Update, CDS, CDS status update 
                               Reply  Wed Sep 5 10:59:23 2018, wgautam, Update, CDS, CDS status update 
                                  Reply  Thu Sep 6 14:21:26 2018, gautam, Update, CDS, ADC replacement in c1lsc expansion chassis CDSoverview.png
                                     Reply  Fri Sep 7 12:35:14 2018, gautam, Update, CDS, ADC replacement in c1lsc expansion chassis Screenshot_from_2018-09-07_12-34-52.png
                                        Reply  Mon Sep 10 12:44:48 2018, Jon, Update, CDS, ADC replacement in c1lsc expansion chassis 
                                     Reply  Thu Sep 20 11:29:04 2018, gautam, Update, CDS, New PCIe fiber housed PCIeFiberSwap.pngPCIeFiberSwap_FBrebooted.png
                                        Reply  Thu Sep 20 16:19:04 2018, gautam, Update, CDS, New PCIe fiber install postponed to tomorrow 
                                           Reply  Fri Sep 21 16:46:38 2018, gautam, Update, CDS, New PCIe fiber installed and routed PCIeFiber.pngIMG_5878.JPG
Message ID: 14140     Entry time: Mon Aug 6 19:49:09 2018     In reply to: 14139     Reply to this: 14142
Author: gautam 
Type: Update 
Category: CDS 
Subject: More CDS woes 

I've left the c1lsc frontend shutdown for now, to see if c1sus and c1ioo can survive without any problems overnight. In parallel, we are going to try and debug the MC1 OSEM Sensor problem - the idea will be to disable the bias voltage to the OSEM LEDs, and see if the readback channels still go below zero, this would be a clear indication that the problem is in the readback transimpedance stage and not the LED. Per the schematic, this can be done by simply disconnecting the two D-sub connectors going to the vacuum flange (this is the configuration in which we usually use the sat box tester kit for example). Attachment #1 shows the current setup at the PD readout board end. The dark DC count (i.e. with the OSEM LEDs off) is ~150 cts, while the nominal level is ~1000 cts, so perhaps this is already indicative of something being broken but let's observe overnight.

Attachment 1: IMG_7106.JPG  1.297 MB  Uploaded Mon Aug 6 21:32:46 2018  | Hide | Hide all
IMG_7106.JPG
ELOG V3.1.3-